OpenStudy (katherinesmith):

Solve the quadratic equation y^2 - 5y = -4 by factoring.

OpenStudy (yamyam70):

First we must equat it to zero :) , what do you think we will do?

OpenStudy (katherinesmith):

y^2 - 5y + 4 = 0

OpenStudy (yamyam70):

Exactly! then we factor out The left hand side :)

OpenStudy (katherinesmith):

can you help me with that? i'm assuming it's trial and error, which i'm really bad at

OpenStudy (katherinesmith):

(y - 1)(y - 4) ?

OpenStudy (yamyam70):

Correct :) But don't forget the zero :)

OpenStudy (katherinesmith):

right. so y = 1 and y = 4

OpenStudy (yamyam70):

thats it :)

OpenStudy (katherinesmith):

thank you! can you help me with another

OpenStudy (yamyam70):

Sure :)

OpenStudy (katherinesmith):

What is the equation of the line which includes points (2, 6) and (0, 0)?

OpenStudy (yamyam70):

Are you given any choices?

OpenStudy (katherinesmith):

y = -3x + 1 y = -3x y = 3x y = 3x + 1

OpenStudy (yamyam70):

I see, We know that in our coordinates, Example (2,6) where, x = 2 , y= 6

OpenStudy (katherinesmith):

okay

OpenStudy (yamyam70):

Then given the following , try solving for x and y :)

OpenStudy (katherinesmith):

wait...what?

OpenStudy (yamyam70):

Substitution, solve for x by substituting the value, which is 2. :)

OpenStudy (anonymous):

y = 1,4

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1}\]

OpenStudy (katherinesmith):

so you want me to substitute 2 in for x in all 4 of the possible answers?

OpenStudy (yamyam70):

yes:)

OpenStudy (anonymous):

what is the equation to find slope?

OpenStudy (yamyam70):

Please @ninab731 start another session.

OpenStudy (katherinesmith):

okay i'm guessing the answer is C

OpenStudy (yamyam70):

lets try out c then :)

OpenStudy (katherinesmith):

i did, i got 6

OpenStudy (yamyam70):

Correct :) so that would be the value for y, now solve for x :)

OpenStudy (katherinesmith):

that gave me -6 = 6

OpenStudy (katherinesmith):

or wait oops 6 = 6

OpenStudy (yamyam70):

when , solving for x, you don't need to substitute x, you just substitute the value of y you got earlier :)

OpenStudy (katherinesmith):

so C

OpenStudy (yamyam70):

for the coordinate (2,6) C is indeed the answer :) I guess you can figure out how to solve (0,0) :)

OpenStudy (katherinesmith):

if i open a new question can you help me? @yamyam70

OpenStudy (yamyam70):

Sure :)